# AppApp Readings

### v1

@@ Abstract
Since there are generally two points ±$\pm{\mathbf{Q}}$ with softening of spin excitations, one should restrict integration in Eq. (3) to the vicinity of one of these points (the presence of two points is accounted for by the factor 1/2 in the right-hand side). The respective region of integration is shown in Fig. S1. In the radial coordinates, it implies
[…]
@@
where $q_{m}(\phi)=\delta/\cos\phi$ for $|\phi|<\arccos(\delta/\Lambda)$ and $q_{m}(\phi)=\Lambda$ otherwise, where $\Lambda$ is a cutoff parameter. In the limit $\delta\rightarrow 0$ (i.e. close to the commensurate phase) we obtain $I=1/2\iint_{q<\Lambda}d^{2}q$, which cancels the factor 1/2 in the right-hand side of Eq. (3) and corresponds to the weight $m^{2}$ of a single mode having $q$ close to $(\pi,\pi)$.
@@
Figure S1: Integration region (shown by colored area) in the sum rule (3)
@@
$\frac{\widetilde{S}^{2}}{2}=T\sum_{i\omega_{n}}\iint\limits_{0}^{\;\;\;\;\Lambda_{\rm uv}}\frac{d^{2}q}{(2\pi)^{2}}\frac{m^{2}}{\rho q^{2}+\chi\omega^{2}+\Delta}$ (A2)
where $\tilde{S}^{2}$ is the square of magnetic moment. Comparing this to the Eq. (3), neglecting $\Delta^{2}$ at $q\sim\Lambda$, and approximating frequency sum by the integral, we find
$\frac{\widetilde{S}^{2}-m^{2}}{2}=T\sum_{i\omega_{n}}\iint\limits_{\Lambda}^{\;\;\;\;\Lambda_{\text{uv}}}\frac{d^{2}q}{(2\pi)^{2}}\frac{m^{2}}{\rho q^{2}+\omega\chi^{2}}\simeq\frac{m^{2}}{4\pi\rho}\int\frac{d\phi}{2\pi}\int\frac{d\omega}{2\pi}\ln\frac{q_{\mathrm{uv}}^{2}(\phi)+(\omega/c)^{2}}{q_{m}^{2}(\phi)+(\omega/c)^{2}}$ (A3)
@@
$\frac{\widetilde{S}^{2}-m^{2}}{2}\simeq\frac{m^{2}c}{4\pi\rho}{\int\frac{d\phi}{2\pi}}(q_{\mathrm{uv}}(\phi)-q_{m}(\phi))\simeq\frac{m^{2}}{4\pi\rho}c(\Lambda_{\mathrm{uv}}-\Lambda)$ (A4)
@@
$\Lambda=\Lambda_{\rm uv}-\frac{\widetilde{S}^{2}-m^{2}}{m^{2}}\frac{2\pi\rho}{c}$ (A5)
Since at $m\rightarrow 0$ we have $\rho\propto\chi\propto m^{2}$, requiring that $\Lambda$ vanishes at $m\rightarrow 0$ (which implies that all degrees of freedom are described qualitatively correct by mean-field theory in that case), we find $\Lambda_{\rm uv}=(\widetilde{S}^{2}/m^{2})\Lambda$ and $\Lambda={2\pi\rho/c}$. Interestingly, the condition $\Lambda\simeq 2\pi\rho/c$ also implies that the second term in Eq. (6) of the main text is of the order of 1, i.e. quantum fluctuations become essential for the chosen momentum cutoff. The quantity $2\pi\rho/c$ is in general temperature- and filling-dependent, but in the considered range $0.950.8<n<$ 0.95 it weakly changes with temperature and doping. Considering the average $\Lambda$, obtained with low- and high-frequency limits of temporal susceptibility, we establish the temperature- and filling-independent cutoff $\Lambda=0.6$.
.2 Mean field approach
At the mean field level for spiral spin density wave magnetic order with general magnetic wave vector ${\bf Q}$ the Hamiltonian (1) can be diagonalized and brought to the form SDW1 ; SDW2 ; SDWIc1 ; SDWIc2 ; SDWIc3 ; SDW3 ; SDW4 ; SDWOur ; OurFirst
$\hat{H}=\sum\limits_{\mathbf{k},\alpha=\pm}E_{\mathbf{k}}^{\alpha}\hat{D}_{\mathbf{k}\alpha}^{+}\hat{D}_{\mathbf{k}\alpha}+NU(m^{2}-\frac{n^{2}}{4}),$ (A6)
@@
$E^{\pm}_{\mathbf{k}}=-\tilde{\mu}+\epsilon^{+}_{\mathbf{k}}\pm\sqrt{(\epsilon_{\mathbf{k}}^{-})^{2}+\Delta^{2}_{\text{el}}},\,\,\,\epsilon^{\pm}_{\mathbf{k}}=\frac{1}{2}(\epsilon_{\mathbf{k}+\mathbf{Q}/2}\pm\epsilon_{\mathbf{k}-\mathbf{Q}/2}),$ (A7)
@@
Figure S2: The dependence of a) local magnetization $m$ and b) normalized magnetic wave vector component $Q_{x}/\pi$ from electron concentration $n$ for different temperatures $T=\frac{1}{\beta}$. Dashed vertical lines correspond to the critical electron concentration $n_{c}$ of the transition between states with AFM and spiral long-range magnetic order.
@@
$W=-T\ln Z=-T\sum\limits_{\mathbf{k\alpha}}\ln\left(1+e^{\frac{-E^{\alpha}_{\mathbf{k}}}{T}}\right)+NU(m^{2}-\frac{n^{2}}{4}).$ (A8)
@@
$1=\frac{U}{2}\sum_{\mathbf{k},\alpha=\pm}\frac{-\alpha f(E^{\alpha}_{\mathbf{k}})}{\sqrt{(\epsilon_{\mathbf{k}}^{-})^{2}+\Delta^{2}_{\text{el}}}},\,\,\,n=\sum_{\mathbf{k},\alpha=\pm}f\left(E_{\mathbf{k}}^{\alpha}\right)$ (A9)
followed by the minimization of Eq. (A8) with respect to the magnetic wave vector $\mathbf{Q}$.
The computed dependence of local magnetization $m$ for states with long-range magnetic order and corresponding magnetic wave vector’s $\mathbf{Q}$ component $Q_{x}$ are presented in Fig. S2. Temperature dependence of both quantities is relatively weak, except for the continuous shift of critical doping level $x_{c}$ of spiral-to-AFM transition to its zero temperature value $x_{c}=0$. There is a gradual shift in the critical electron concentration, corresponding to the phase transition between spiral and AFM magnetic orders, towards half-filling as the temperature decreases. The obtained order and incommensurability parameters coincide with the results presented in Ref. BonettiThesis, of the paper.
.3 Calculation of spin susceptibility and spin current correlation functions
From Eq. (2) of the paper, it follows that the spatial spin stiffness can be directly obtained as $\chi_{\rm{op},\omega_{n}}={m^{2}}/({\bar{\chi}^{yy}_{\mathbf{Q},\omega_{n}}\omega_{n}^{2}})$. For the computation of the nonuniform dynamical spin susceptibility in the local coordinate frame, $\bar{\chi}^{\alpha,\beta}_{\mathbf{q}}$, we employ the random phase approximation (RPA), following the expression
$\bar{\chi}_{\mathbf{q}}=(\hat{I}-\bar{\chi}^{(0)}_{\mathbf{q}}\hat{U})^{-1}\bar{\chi}^{(0)}_{\mathbf{q}},$ (A10)
@@
$\hat{U}=\left(\begin{matrix}0&0&0&-U\\ 0&U&0&0\\ 0&0&U&0\\ -U&0&0&0\end{matrix}\right),$ (A11)
@@
$\chi^{(0)\;\alpha\beta}_{q,\omega_{n}}=-\frac{T}{4}\sum\limits_{k}\mathrm{Tr}[\sigma^{\alpha}G_{k}\sigma^{\beta}G_{k+q}]$ (A12)
@@
$G_{\mathbf{k},\nu}^{\sigma\sigma^{\prime}}=\begin{pmatrix}i\nu+\mu-Un_{\downarrow}-\epsilon^{+}_{\mathbf{k}}&{\rm i}\epsilon^{-}_{\mathbf{k}}\\ -{\rm i}\epsilon^{-}_{\mathbf{k}}&i\nu+\mu-Un_{\uparrow}-\epsilon^{+}_{\mathbf{k}}\end{pmatrix}^{-1}$ (A13)
[…]
The susceptibility $\bar{\chi}_{\mathbf{q}}$ obtained from Eq. (A10) serves as a key quantity for evaluating the dynamical spin stiffness $\chi_{\text{op},\omega_{n}}$ and analyzing the stability of mean-field solutions against the emergence of inhomogeneous phases by examining the spectrum of $\bar{\chi}_{\mathbf{q}}$ throughout the entire Brillouin zone.
The spatial spin stiffness of the out-of-plane $\rho_{\text{op},\mu}$ and in-plane $\rho_{\text{ip},\mu}$ modes with $\mu=x,y$ are related to the full gauge kernel function $M^{xx}_{\mathbf{q},\omega_{n}}$ in the mean-field approximation can be expressed as OurSecond ; BonettiThesis ; Bonetti ; BonettiMetzner
$\rho_{\text{op},n}$ $=2\left({K}_{0,nn}^{(0),xx}+{K}_{0,nn}^{d}\right),$ (A14)
$\rho_{\text{ip},n}$ $={K}_{0,nn}^{(0),yy}+{K}_{0,nn}^{d},$ (A15)
where $K^{d}_{\mu\nu}$ accounts for the diamagnetic contribution, and ${K}_{0,nn}^{(0),xx}$ is a bubble contribution to the correlation function $\langle\langle\hat{J}_{\mathbf{q},{\mu}}^{\alpha}|\hat{J}_{-\mathbf{q},{\nu}}^{\beta}\rangle\rangle$ of spin currents $\hat{J}^{\alpha}_{\mathbf{q},\mu}$, $\alpha=x,y,z,0$ ($\alpha=0$ corresponds to the usual charge current operator), $\mu=x,y$. In $\rho_{in,n}$ we have also neglected the contribution $\Delta\rho_{n}$ discussed previously in Ref. OurSecond, , originating from mixing of different susceptibility components, as it was found to be numerically small. Similarly to the equation (A12) one can express ${K}_{0,nn}^{(0),xx}$ and $K^{d}_{0,nn}$ in terms of single-particle Green functions
${K}_{0,nn}^{(0),xx}$ $=\sum\limits_{s=\pm}\left[\bar{K}^{(0),xx}-is(\bar{K}^{(0),xz}-\bar{K}^{(0),zx})+\bar{K}^{(0),zz}\right]_{q,nn,ss},$ (A16)
$K_{q,\mu\nu}^{(0),yy}$ $=\bar{K}_{q,\mu\nu,++}^{(0),yy}+\bar{K}_{q,\mu\nu,--}^{(0),00}+\bar{K}_{q,\mu\nu,+-}^{(0),y0}+\bar{K}_{q,\mu\nu,-+}^{(0),0y},$ (A17)
$\bar{K}^{0;\alpha\beta}_{q;\mu\rho,ss^{\prime}}$ $=T\sum_{k}T^{\mu,s}_{{\mathbf{k},\nu},{\mathbf{q}}}\mathrm{Tr}\left[\sigma^{\alpha}G_{k}\sigma^{\beta}G_{k+q}\right]T^{\rho,s^{\prime}}_{{\mathbf{k}},{\mathbf{q}}},$ (A18)
@@
$T^{\mu\pm}_{{\mathbf{k}},{\mathbf{q}}}$ $=T^{\mu}_{{\mathbf{k}}+s{\mathbf{Q}}/2,{\mathbf{q}}}\,\,\,$ $(\alpha,\beta=x,z),$ (A19)
@@
$\bar{K}_{\mu\nu,s}^{d,ab}=-\delta_{ab}(1-\delta_{\mu 0})(1-\delta_{\nu 0})\sum_{k}T^{\mu\nu,s}_{\mathbf{k}}\mathrm{Tr}\left[\sigma^{a}G_{k}\right].$ (A20)
.4 Spatial spin stiffnesses
@@
For completeness, we show the dependence of spatial spin stiffnesses calculated via equations (A14) and (A15) on electron concentration, see Fig. S4. The temperature dependence of spatial spin stiffness is different for the cases of out-of-plane and in-plane responses. As temperature decreases the dependence of the in-plane spatial stiffness $\rho_{2,xx/yy}$ becomes sharper in the vicinity of half-filling. With $T\to 0$ it converges at all finite doping levels to the value nearly independent of electron concentration and develops a jump at the half-filling.
@@
Figure S3: The dependence of spatial spin stiffness on electron concentration $n$ for various temperatures $T=\frac{1}{\beta}$ for the out-of-plane (left) and in-plane (right) excitations.
.5 Frequency dependence of the in-plane temporal spin stiffness

### v3

@@ 1 Connection to the Nonlinear Sigma Model
Since there are generally two points $\mathbf{q}=\pm{\mathbf{Q}}$ with softening of spin excitations, one should restrict integration in Eq. (2) of the main text to the vicinity of one of these points (the presence of two points is accounted for by the factor 1/2 in the right-hand side). The respective region of integration is shown in Fig. 1. In the radial coordinates, it implies
$I=\iint d^{2}q=\int\limits_{-\pi}^{\pi}d\phi\int\limits_{0}^{q_{m}(\phi)}qdq,$ (S11)
@@
where $q_{m}(\phi)=\delta/\cos\phi$ for $|\phi|<\arccos(\delta/\Lambda)$ and $q_{m}(\phi)=\Lambda$ otherwise, where $\Lambda$ is a cutoff parameter. In the limit $\delta\rightarrow 0$ (i.e. close to the commensurate phase) we obtain $I=1/2\iint_{q<\Lambda}d^{2}q$, which cancels the factor 1/2 in the right-hand side of Eq. (2) of the main text and corresponds to the weight $m^{2}$ of a single mode with $q$ close to $(\pi,\pi)$.
@@
Figure 1: Integration region (shown by colored area) in the sum rule of Eq. (2) of the main text
@@
$\frac{\widetilde{S}^{2}}{2}=T\sum_{i\omega_{n}}\iint\limits_{0}^{\;\;\;\;\Lambda_{\rm uv}}\frac{d^{2}q}{(2\pi)^{2}}\frac{m^{2}}{\rho q^{2}+\chi\omega^{2}+\Delta}$ (S12)
where $\tilde{S}^{2}$ is the square of magnetic moment. Comparing this to the Eq. (2) of the main text, neglecting $\Delta^{2}$ at $q\sim\Lambda$, and approximating frequency sum by the integral, we find
$\frac{\widetilde{S}^{2}-m^{2}}{2}=T\sum_{i\omega_{n}}\iint\limits_{\Lambda}^{\;\;\;\;\Lambda_{\text{uv}}}\frac{d^{2}q}{(2\pi)^{2}}\frac{m^{2}}{\rho q^{2}+\omega\chi^{2}}\simeq\frac{m^{2}}{4\pi\rho}\int\frac{d\phi}{2\pi}\int\frac{d\omega}{2\pi}\ln\frac{q_{\mathrm{uv}}^{2}(\phi)+(\omega/c)^{2}}{q_{m}^{2}(\phi)+(\omega/c)^{2}}$ (S13)
@@
$\frac{\widetilde{S}^{2}-m^{2}}{2}\simeq\frac{m^{2}c}{4\pi\rho}{\int\frac{d\phi}{2\pi}}(q_{\mathrm{uv}}(\phi)-q_{m}(\phi))\simeq\frac{m^{2}}{4\pi\rho}c(\Lambda_{\mathrm{uv}}-\Lambda)$ (S14)
@@
$\Lambda=\Lambda_{\rm uv}-\frac{\widetilde{S}^{2}-m^{2}}{m^{2}}\frac{2\pi\rho}{c}$ (S15)
Since at $m\rightarrow 0$ we have $\rho\propto\chi\propto m^{2}$, requiring that $\Lambda$ vanishes at $m\rightarrow 0$ (which implies that all degrees of freedom are described qualitatively correct by mean-field theory in that case), we find $\Lambda_{\rm uv}=(\widetilde{S}^{2}/m^{2})\Lambda$ and $\Lambda={2\pi\rho/c}$. Interestingly, the condition $\Lambda\simeq 2\pi\rho/c$ also implies that the second term in Eq. (5) of the main text is of the order of 1, i.e. quantum fluctuations become essential for the chosen momentum cutoff. The quantity $2\pi\rho/c$ is in general temperature- and filling-dependent, but in the considered range $0.8<n<0.95$ it weakly changes with temperature and doping. Considering the average $\Lambda$, obtained with low- and high-frequency limits of temporal susceptibility, we establish the temperature- and filling-independent cutoff $\Lambda=0.6$.
3 Mean field approach
At the mean field level for spiral spin density wave magnetic order with general magnetic wave vector $\bf Q$} the Hamiltonian of Eq. (8) of the main text can be diagonalized and brought to the form SDW1 ; SDW2 ; SDWIc1 ; SDWIc2 ; SDWIc3 ; SDW3 ; SDW4 ; SDWOur ; OurFirst
$\hat{H}=\sum\limits_{\mathbf{k},\alpha=\pm}E_{\mathbf{k}}^{\alpha}\hat{D}_{\mathbf{k}\alpha}^{+}\hat{D}_{\mathbf{k}\alpha}+NU(m^{2}-\frac{n^{2}}{4}),$ (S16)
@@ 3 Mean field approach
$E^{\pm}_{\mathbf{k}}=-\tilde{\mu}+\epsilon^{+}_{\mathbf{k}}\pm\sqrt{(\epsilon_{\mathbf{k}}^{-})^{2}+\Delta^{2}_{\text{el}}},\,\,\,\epsilon^{\pm}_{\mathbf{k}}=\frac{1}{2}(\epsilon_{\mathbf{k}+\mathbf{Q}/2}\pm\epsilon_{\mathbf{k}-\mathbf{Q}/2}),$ (S17)
@@
Figure 2: The dependence of a) local magnetization $m$ and b) normalized magnetic wave vector component $Q_{x}/\pi$ from electron concentration $n$ for different temperatures $T=\frac{1}{\beta}$. Dashed vertical lines correspond to the critical electron concentration $n_{c}$ of the transition between states with AFM and spiral long-range magnetic order.
@@
$W=-T\ln Z=-T\sum\limits_{\mathbf{k\alpha}}\ln\left(1+e^{\frac{-E^{\alpha}_{\mathbf{k}}}{T}}\right)+NU(m^{2}-\frac{n^{2}}{4}).$ (S18)
@@
$1=\frac{U}{2}\sum_{\mathbf{k},\alpha=\pm}\frac{-\alpha f(E^{\alpha}_{\mathbf{k}})}{\sqrt{(\epsilon_{\mathbf{k}}^{-})^{2}+\Delta^{2}_{\text{el}}}},\,\,\,n=\sum_{\mathbf{k},\alpha=\pm}f\left(E_{\mathbf{k}}^{\alpha}\right)$ (S19)
followed by the minimization of Eq. (S18) with respect to the magnetic wave vector $\mathbf{Q}$.
The computed dependence of local magnetization $m$ for states with long-range magnetic order and corresponding magnetic wave vector’s $\mathbf{Q}$ component $Q_{x}$ are presented in Fig. 2. Temperature dependence of both quantities is relatively weak, except for the continuous shift of critical doping level $x_{c}$ of spiral-to-AFM transition to its zero temperature value $x_{c}=0$. There is a gradual shift in the critical electron concentration, corresponding to the phase transition between spiral and AFM magnetic orders, towards half-filling as the temperature decreases. The obtained order and incommensurability parameters coincide with the results presented in Ref. BonettiThesis, of the main text.
4 Calculation of spin susceptibility and spin current correlation functions
From Eq. (1) of the main text, it follows that the spatial spin stiffness can be directly obtained as $\chi_{\rm{op},\omega_{n}}={m^{2}}/({\bar{\chi}^{yy}_{\mathbf{Q},\omega_{n}}\omega_{n}^{2}})$. For the computation of the nonuniform dynamical spin susceptibility in the local coordinate frame, $\bar{\chi}^{\alpha,\beta}_{\mathbf{q}}$, we employ the random phase approximation (RPA), following the expression
$\bar{\chi}_{\mathbf{q}}=(\hat{I}-\bar{\chi}^{(0)}_{\mathbf{q}}\hat{U})^{-1}\bar{\chi}^{(0)}_{\mathbf{q}}$, (S20)
@@ 4 Calculation of spin susceptibility and spin current correlation functions
$\hat{U}=\left(\begin{matrix}0&0&0&-U\\ 0&U&0&0\\ 0&0&U&0\\ -U&0&0&0\end{matrix}\right),$ (S21)
@@
$\chi^{(0)\;\alpha\beta}_{q,\omega_{n}}=-\frac{T}{4}\sum\limits_{k}\mathrm{Tr}[\sigma^{\alpha}G_{k}\sigma^{\beta}G_{k+q}]$ (S22)
@@
$G_{\mathbf{k},\nu}^{\sigma\sigma^{\prime}}=\begin{pmatrix}i\nu+\mu-Un_{\downarrow}-\epsilon^{+}_{\mathbf{k}}&{\rm i}\epsilon^{-}_{\mathbf{k}}\\ -{\rm i}\epsilon^{-}_{\mathbf{k}}&i\nu+\mu-Un_{\uparrow}-\epsilon^{+}_{\mathbf{k}}\end{pmatrix}^{-1}$ (S23)
being the one-particle Green functions in local coordinate frame; $\epsilon_{\mathbf{k}}^{\pm}$ is defined according to (S17).
The susceptibility $\bar{\chi}_{\mathbf{q}}$ obtained from Eq. (S20) serves as a key quantity for evaluating the dynamical spin stiffness $\chi_{\text{op},\omega_{n}}$ and analyzing the stability of mean-field solutions against the emergence of inhomogeneous phases by examining the spectrum of $\bar{\chi}_{\mathbf{q}}$ throughout the entire Brillouin zone.
The spatial spin stiffness of the out-of-plane $\rho_{\text{op},\mu}$ and in-plane $\rho_{\text{ip},\mu}$ modes with $\mu=x,y$ are related to the full gauge kernel function $M^{xx}_{\mathbf{q},\omega_{n}}$ in the mean-field approximation can be expressed as (Refs. OurSecond ; BonettiThesis ; Bonetti ; BonettiMetzner of the main text)
$\rho_{\text{op},n}$ $=2\left({K}_{0,nn}^{(0),xx}+{K}_{0,nn}^{d}\right),$ (S24)
$\rho_{\text{ip},n}$ $={K}_{0,nn}^{(0),yy}+{K}_{0,nn}^{d},$ (S25)
where $K^{d}_{\mu\nu}$ accounts for the diamagnetic contribution, and ${K}_{0,nn}^{(0),xx}$ is a bubble contribution to the correlation function $\langle\langle\hat{J}_{\mathbf{q},{\mu}}^{\alpha}|\hat{J}_{-\mathbf{q},{\nu}}^{\beta}\rangle\rangle$ of spin currents $\hat{J}^{\alpha}_{\mathbf{q},\mu}$, $\alpha=x,y,z,0$ ($\alpha=0$ corresponds to the usual charge current operator), $\mu=x,y$. In $\rho_{in,n}$ we have also neglected the contribution $\Delta\rho_{n}$ discussed previously in Ref. OurSecond, of the main text, originating from mixing of different susceptibility components, as it was found to be numerically small. Similarly to the equation (S22) one can express ${K}_{0,nn}^{(0),xx}$ and $K^{d}_{0,nn}$ in terms of single-particle Green functions
${K}_{0,nn}^{(0),xx}$ $=\sum\limits_{s=\pm}\left[\bar{K}^{(0),xx}-is(\bar{K}^{(0),xz}-\bar{K}^{(0),zx})+\bar{K}^{(0),zz}\right]_{q,nn,ss},$ (S26)
$K_{q,\mu\nu}^{(0),yy}$ $=\bar{K}_{q,\mu\nu,++}^{(0),yy}+\bar{K}_{q,\mu\nu,--}^{(0),00}+\bar{K}_{q,\mu\nu,+-}^{(0),y0}+\bar{K}_{q,\mu\nu,-+}^{(0),0y},$ (S27)
$\bar{K}^{0;\alpha\beta}_{q;\mu\rho,ss^{\prime}}$ $=T\sum_{k}T^{\mu,s}_{{\mathbf{k},\nu},{\mathbf{q}}}\mathrm{Tr}\left[\sigma^{\alpha}G_{k}\sigma^{\beta}G_{k+q}\right]T^{\rho,s^{\prime}}_{{\mathbf{k}},{\mathbf{q}}},$ (S28)
@@
$T^{\mu\pm}_{{\mathbf{k}},{\mathbf{q}}}$ $=T^{\mu}_{{\mathbf{k}}+s{\mathbf{Q}}/2,{\mathbf{q}}}\,\,\,$ $(\alpha,\beta=x,z),$ (S29)
@@
$\bar{K}_{\mu\nu,s}^{d,ab}=-\delta_{ab}(1-\delta_{\mu 0})(1-\delta_{\nu 0})\sum_{k}T^{\mu\nu,s}_{\mathbf{k}}\mathrm{Tr}\left[\sigma^{a}G_{k}\right].$ (S30)
5 Spatial spin stiffnesses
@@
For completeness, we show the dependence of spatial spin stiffnesses calculated via equations (S24) and (S25) on electron concentration, see Fig. 3. The temperature dependence of spatial spin stiffness is different for the cases of out-of-plane and in-plane responses. As temperature decreases the dependence of the in-plane spatial stiffness $\rho_{2,xx/yy}$ becomes sharper in the vicinity of half-filling. With $T\to 0$ it converges at all finite doping levels to the value nearly independent of electron concentration and develops a jump at the half-filling.
@@
Figure 3: The dependence of spatial spin stiffness on electron concentration $n$ for various temperatures $T=\frac{1}{\beta}$ for the out-of-plane (left) and in-plane (right) excitations.
6 Frequency dependence of the in-plane temporal spin stiffness